the sum of two fractions is 1. if one fractions is 3/8 greater than the other, what are the two fractions? If I take 8/8 - 3/8 = 5/8 and that equals 1 but how is that 3/8 greater than the other? It isn't right? Please help me show my work.
Thank you,
OP

Answer by lwsshak3(11628)

he sum of two fractions is 1. if one fractions is 3/8 greater than the other, what are the two fractions?
let x=one of the fractions
x+3/8=the other fraction
x+x+3/8=1
2x=1/38=5/8
x=5/16
x+3/8=5/16+6/16=11/16
The two fractions are: 5/16 and 11/16